A metal bowl with a weight of 1.75 N is placed in a larger kitchen container filled with corn oil. How much corn oil must the bowl displace in order to float? For reference, the mass density of corn oil is about 900 g/liter and its weight density is about 8.82 N/liter. Please give your answer in liters.
1
Expert's answer
2019-04-12T09:39:56-0400

According to Archimedes law 8.82 (N / l) * V = 1.75 (N), therefore V = 1.75 / 8.82 = 0.2 (l), so V = 0.2 l.